A truck with a mass of 2300kg is moving north at 50km/hr when it collides with a car of mass 1100kg moving east at 75km/hr. Assuming a perfectly inelastic collision, find the speed and angle (relative to east) of the wreckage.
Answer #1

Mv_{north}=(m+M)V_{north}

V_{north}=\frac{2300*50}{(1100+2300)}=33.82\ km/hr

mv_{east}=(m+M)V_{east}

V_{east}=\frac{1100*75}{(1100+2300)}=24.26\ km/hr

Speed=\sqrt{24.26^{2}+33.82^{2}}=41.62\ km/hr

\Theta = tan^{-1}\frac{33.82}{24.26}=54.35^{o}